OpenStudy (anonymous):

First derivative of y= 10^cos3x

OpenStudy (anonymous):

\[y= 10^{\cos3x}\]

OpenStudy (kainui):

Show your best guess and we'll help you. =)

OpenStudy (anonymous):

my answer was - 30 cos 3x

OpenStudy (anonymous):

it was wrong..

OpenStudy (anonymous):

wonr, its -30sin3x.

OpenStudy (anonymous):

yep it's wrong

OpenStudy (kainui):

You can rewrite it like this: \[y=e^{\ln10*\cos3x}\] I think it makes it easier to think about when it's like this.

OpenStudy (anonymous):

the derivative of \[a ^{u} = u \prime \times a ^{u} \times lna\]

OpenStudy (anonymous):

@silverxx can u try to solve it ?

OpenStudy (anonymous):

-3 sin3x (10) ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

really have no idea ;(

OpenStudy (kainui):

You need to use the chain rule here.

OpenStudy (anonymous):

listen the derivative of a^u=u′×a^u×lna u have 10^cos3x u′=-3sin3x a^u=10^cos3x lna=ln10 so the derivative of 10^cos3x=-3sin3x*10^cos3x*ln10

OpenStudy (anonymous):

is it.. cos 3x [ (10) (-3sin3x)]?

OpenStudy (anonymous):

it's \[-3\sin3x \times10^{\cos3x}\times \ln 10\]

OpenStudy (anonymous):

@silverxx

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

woah. tsk. Thank you so much!! again.

OpenStudy (anonymous):

Anytime. :D